### v2

# Dynamic phase transition into a mixed-CDW state in 1T-TaS2 via a thermal quench

Dynamic phase transition into a mixed-CDW state in 1T-TaS2 via a thermal quench
A. de la Torre a.delatorreduran@northeastern.edu Quantum Materials and Sensing Institute, Northeastern University, Burlington, Massachusetts, 01803, USA Department of Physics, Northeastern University, Boston, Massachusetts, 02115, USA Department of Physics, Brown University, Providence, Rhode Island 02912, United States    Q. Wang Department of Physics, Brown University, Providence, Rhode Island 02912, United States    B. Campbell Department of Physics and Astronomy, University of New Hampshire, Durham, NH 03824.    J. V. Riffle Department of Physics and Astronomy, University of New Hampshire, Durham, NH 03824.    D. Balasundaram Department of Physics and Astronomy, George Mason University, Fairfax, VA, 22030 USA Quantum Science and Engineering Center, George Mason University, Fairfax, VA, 22030 USA    P.M.Vora Department of Physics and Astronomy, George Mason University, Fairfax, VA, 22030 USA Quantum Science and Engineering Center, George Mason University, Fairfax, VA, 22030 USA    J.P.C. Ruff Cornell High Energy Synchrotron Source, Cornell University, Ithaca, NY, 14853, USA    S. M. Hollen Department of Physics and Astronomy, University of New Hampshire, Durham, NH 03824.    K. W. Plumb kemp˙plumb@brown.edu Department of Physics, Brown University, Providence, Rhode Island 02912, United States
(July 12, 2024)
Ultrafast light-matter interaction has emerged as a new mechanism to exert control over the macroscopic properties of quantum materials toward novel functionality. To date, technological applications of these non-thermal phases are limited by their ultrashort lifetimes and low-ordering temperatures. Among the most studied photoinduced metastable phases for their technological promise is the hidden metallic charge density wave (H-CDW) in the model correlated CDW compound $1T$-TaS2. Despite active study and engineering, the nature of the photoinduced H-CDW remains the subject of debate and is only accessible at cryogenic temperatures. Here, we stabilize the H-CDW phase at thermal equilibrium up to near-room temperature by accessing an intermediate mixed CDW order regime via thermal quenching. Using x-ray high dynamic range reciprocal space mapping (HDRM) and scanning tunneling spectroscopy (STS), we reveal the coexistence of commensurate (C) CDW and H-CDW domains below 180 K during cooling and below 210 K during warming. Our findings show that each order parameter breaks basal plane mirror symmetry with different chiral orientations and induces out-of-plane unit cell tripling in the H-CDW phase. Despite metallic domain walls and a finite density of states at zero bias observed via STS, bulk resistance remains insulating due to CDW stacking disorder. This study establishes the H-CDW as a thermally stable phase and introduces a new mechanism for switchable metallic behavior in thin flakes of $1T$-TaS2 and similar materials with competing order phases.
I Introduction
The correlated van der Waal material $1T$-TaS2 displays a series of first-order phase transitions associated with the onset of CDW orders with different degrees of commensurability Thompson et al. (1971); Wilson et al. (1975); Fazekas and Tosatti (1979). First, at $T=550$ K, a mirror-symmetry preserving incommensurate (I) CDW onsets from a metallic high-temperature state Sipos et al. (2008). At $T_{\mathrm{I-NC}}$$\approx 350$ K, $1T$-TaS2 enters into a nearly commensurate (NC) CDW phase. The atomic rearrangement associated with the incommensurate NC-CDW (in-plane ordering wavevector $q_{||}\!=\!0.284$ r.l.u.) leads to disconnected regions displaying a $\sqrt{13}\times\sqrt{13}$ structural reconstruction in which twelve Ta atoms move towards a central one forming a star-of-David (SOD) pattern Spijkerman et al. (1997). The NC-CDW breaks the mirror plane symmetries of the lattice basal plane and spontaneously selects one of two chiral domains rotated from the $a$-axis by $\phi\approx 12^{\circ}$. Then, at $T_{\mathrm{NC-C}}\approx 175$ K in cooling and $T_{\mathrm{NC-C}}\approx 225$ K in warming, a 3D CDW order emerges from the NC-CDW phase. This low-temperature phase is commensurate (C) with the atomic lattice in-plane ($\phi\!=\!13.9^{\circ}$ and $q_{||}\!=\!0.277$ r.l.u.) with the in-plane C-CDW atomic reconstruction inheriting the chiral structure from the NC-CDW. On the other hand, in the out-of-plane direction, the C-CDW displays a large stacking disorder between dimerized CDW planes Thompson et al. (1971); Wilson et al. (1975).
In general, experimental probes spanning a large range of lengthscales find that $1T$-TaS2 samples lack domain twinning that is expected across a first-order phase transition Zong et al. (2018); Fichera et al. (2020); Luo et al. (2021); Yang et al. (2022); Campbell et al. (2024). This reflects the kinetics of the first cooling across the only first-order phase transition in $1T$-TaS2 that breaks in-plane mirror symmetries, $T_{\mathrm{I-NC}}$ Sept and Tuszyński (1994); Binder (1987). During synthesis, the $1T$ phase is stabilized by quenching to $T\approx 273$ K across the polytype transition, $T_{\rm{1T-2H}}\!=\!600$ K Thompson et al. (1971). During this initial quench, the cooling rate across $T_{\mathrm{I-NC}}$ is effectively slow when compared to the intrinsic timescales of the electronic ($\approx$ fs) and atomic ($\approx$ ps) degrees of freedom and enables the nucleation of large chiral CDW domains with lateral sizes comparable to the correlation length of the order parameter parameter Oxtoby (1998). The chiral domain structure of the C-CDW in $1T$-TaS2 is stable to cooling through $T_{\mathrm{NC-C}}$ Zong et al. (2018); Fichera et al. (2020); Luo et al. (2021); Yang et al. (2022); Campbell et al. (2024) with suggestions that strain fields and defects can pin a single chirality Zong et al. (2018); Salzmann et al. (2023a). Still, chirality reversal of single domain samples has been demonstrated by repeated cleaving of a bulk sample Fichera et al. (2020) or by cycling through $T_{\mathrm{I-NC}}$ Zhao et al. (2023). On the other hand, multi-chiral domain samples can be engineered in thin exfoliated flakes by heating and quenching above $T_{\rm{1T-2H}}$ Sung et al. (2022); Husremović et al. (2023) or upon irradiation with 800 nm laser pulses with fluence $F>7$ mJ/cm2 Zong et al. (2018). Once created, chiral domain walls are long-lived at low temperatures, but a dominant chiral domain of dimensions comparable to the sample size can be restored by heating above $T_{\mathrm{I-NC}}$ and adiabatically cooling through the mirror symmetry-breaking IC-NC transition Sung et al. (2022); Husremović et al. (2023); Zong et al. (2018).Thus, $T_{\mathrm{I-NC}}$ emerges as the relevant energy scale for chiral CDW domain nucleation in $1T$-TaS2 as the initial mirror-symmetry breaking transition in the CDW phase diagram Thompson et al. (1971). Moreover, x-ray pair distribution function has revealed that the in-plane Ta average displacement from the mean crystallographic position, a measure of the CDW formation, remains constant across the NC-C CDW transition. At the same time, it dramatically changes across $T_{\mathrm{I-NC}}$ Bozin et al. (2023).
[…]
An additional low-temperature metastable metallic CDW phase, the so-called hidden H-CDW phase, has been accessed in $1T$-TaS2 from the insulating C-CDW state via a single femtosecond laser pulse of fluence larger than a critical value ($F_{c}\!>\!1$ mJ/cm2) Stojchevska et al. (2014); Stahl et al. (2020); Vaskivskyi et al. (2015) or by exciting $1T$-TaS2 with a ns voltage pulse Hollander et al. (2015); Ma et al. (2016); Cho et al. (2016); Venturini et al. (2022); Tsen et al. (2015). The H-CDW phase locally resembles the SOD arrangement but breaks the large C-CDW domains into many nanoscale domains and has a metallic density of states Gerasimenko et al. (2019); Ma et al. (2016); Maklar et al. (2023) These domains maintain the chirality of the C-CDW Stahl et al. (2020) and are related by a relative translation of the SOD pattern by a finite number of atomic lattice vectors Ma et al. (2016); Gerasimenko et al. (2019); Ravnik et al. (2021). Moreover, slight differences between the displacement of the Ta atoms between the H-CDW and the C-CDW distinguish the two. The H-CDW phases result in CDW Bragg peaks rotated by $\phi\!=\!11.9^{\circ}$ from the $a^{*}$-axis with a larger basal plane ordering wavevector, $q_{||}\!=\!0.284$ r.l.u.. The CDW plane stacking also rearranges towards a $l\!=\!1/3$ periodicity along the c-axis Stahl et al. (2020). We note that a similar metallic mosaic state can also be induced by cooling thin exfoliated samples ($d<24$ nm) across $T_{\mathrm{I-NC}}$ with a cooling rate faster than $0.2$ K/s Yoshida et al. (2015) or by intrinsic strain and defect effects Mutka et al. (1981); Zhang et al. (2022); Salzmann et al. (2023b).
Despite much study, the origin of the H-CDW state remains controversial, with photodoping Sun et al. (2018); Gerasimenko et al. (2019), incoherent thermal effects Ma et al. (2016); Gao et al. (2022) or electron-phonon relaxation bottlenecks Maklar et al. (2023); Mante et al. (2022) all possible explanations. Nonetheless, the H-CDW has attracted much interest, not only as an example of the non-thermal control of quantum materials de la Torre et al. (2021) but also as a platform for new functional devices Mihailovic et al. (2021); Mraz et al. (2022); Devidas et al. (2024). However, the H-CDW has only been stabilized for ultrashort periods (${}100\mu$s) at temperatures above $T\!=\!50$ K Stojchevska et al. (2014); Ravnik et al. (2018); Maklar et al. (2023) preventing its deployment in efficient and fast memory devices towards novel technological applications Basov et al. (2017).
Here, we demonstrate that the H-CDW phase exists at thermal equilibrium in $1T$-TaS2 and extend its thermal stability to near-room temperature by accessing a new long-lived mixed-CDW state via a reversible thermal quench across $T_{\mathrm{I-NC}}$. We deploy x-ray high dynamic range reciprocal space mapping (HDRM) to unequivocally demonstrate the co-existence of the C-CDW and H-CDW states in thermally quenched samples. Each CDW order parameter is linked to a mirror-symmetry breaking chiral domain. The mixed-CDW state emerges from a multi-domain NC-CDW state below $T^{CD}=180$ K in cooling and $T^{WU}=210$ K in warming. The mixed-CDW remains stable up to $T_{\mathrm{I-NC}}$. Scanning tunneling spectroscopy (STS) shows that the mixed-CDW is characterized by a semi-metallic in-plane density of states. At the same time, bulk resistivity remains insulating due to out-of-plane CDW stacking disorder in mixed CDW bulk samples. Our results demonstrate a new state in the phase diagram of $1T$-TaS2, enabling a new mechanism to engineer a semimetallic density of states at near room temperature.
[…]
II Results
Accessing a novel mixed-CDW state by quenching through the mirror symmetry breaking transition. In Figure 1 a, we show a characteristic [HK0] plane reciprocal space map for the C-CDW phase of $1T$-TaS2, collected after a standard cooling cycle. Red circular markers highlight the CDW reflections at $q^{C}_{1}\!=\!(\sigma^{C}_{1},\sigma^{C}_{2},l)$ and $q^{TH}_{2}\!=\!(-\sigma^{C}_{2},\sigma^{C}_{1}+\sigma^{C}_{2}+,l)$ with $\sigma^{C}_{1}\!=\!3/13$ r.l.u and $\sigma^{C}_{2}\!=\!1/13$ r.l.u Thompson et al. (1971); Wilson et al. (1975), corresponding to a single C-CDW chiral domain. Additional non-indexable reflections in the reciprocal space map are attributed to a second small crystallite in our sample. Identical reciprocal space maps were obtained across many cooling cycles and samples. However, we found that the C-CDW domain structure depends on the cooling rate, and a fast thermal quench across $T_{\mathrm{I-NC}}$ can consistently produce a multi-domain CDW. Fig. 1 b shows the [HK0] reciprocal space map after an in-situ quench in our bulk sample across $T_{\mathrm{I-NC}}$ down to $T\!=\!140$ K at a rate of $0.2$ K/ms near $T_{\mathrm{I-NC}}$ (See Methods for a description of the in-situ quench method and calculation of quenching rates using heat diffusion calculations). We found that thermal quenching consistently resulted in a new set of superlattice reflections of opposite chirality to the original set of satellite peaks ($\phi\!=\!-12.3^{\circ}$) [Fig. 1 b]. The superior reciprocal space resolution of synchrotron single crystal HDRM enables us to resolve differences in the wavevector of the two sets of CDW Bragg peaks with different chirality. The in-plane wavevector magnitude $|q_{||}|\!=\!0.284(3)~{}\AA^{-1}$ of the quenched-induced satellite peaks is larger than that of the C-CDW ($|q_{||}|\!=\!0.277(3)~{}\AA^{-1}$) and agrees with that of the light-stabilized H-CDW state Stahl et al. (2020). This is shown in Fig. 1 c, where we display the integrated intensity across the C-CDW and H-CDW Bragg peaks (dashed lines in Fig. 1 a and b) as a function of $|q_{||}|$. Strikingly, each order parameter preferentially selects a chirality orientation, with the C-CDW domains of the mixed-CDW phase maintaining that of the as-grown samples.
We also find that the c-axis periodicity differs between the C-CDW and H-CDW-like domains. Figure 1 d and e show [H0L] reciprocal space maps centered at H = 1 r.l.u and L = 0 r.l.u before and after the thermal quench, respectively. The as-grown sample displays a broad two-peak structure, with the center of mass at integer values of L. These short-range out-of-plane correlations are characteristic of the dimerization of CDW layers in the low-temperature C-CDW phase (Fig. 1 g) Tanda et al. (1984). On the other hand, the H-CDW-like satellite peaks shift towards $l\!=\!\pm 1/3$ indicating a collapse of the interlayer dimerization towards a triple layer CDW stacking (Fig. 1 h) as was observed for the light-induced H-CDW Stahl et al. (2020). The H-CDW-like satellite peaks are sharper than the C-CDW ones, reflecting a longer correlation length for the quench-induced domains $\xi\!=\!6(6)$ nm (Fig. 1 f). Our HDRM data in quenched samples demonstrates the coexistence of the C-CDW with a second CDW phase with the same order parameter as the H-CDW phase at much higher temperatures ($T=140$ K).
Critical temperature and metastability of the quenched mixed-CDW state. We performed a set of in-situ thermal cycles to directly demonstrate that (i) $T_{\mathrm{I-NC}}$ is the relevant temperature scale to access the mixed state and (ii) the transition is reversible. Consecutive thermal quenches were performed at different quench rates and starting temperatures, as illustrated in Figure 2 a (see Methods and Extended data). We found that after quenching from above $T_{\mathrm{I-NC}}$, the mixed CDW state is thermally stable and insensitive to warming-cooling cycles if temperature kept less than $T_{\mathrm{I-NC}}$ [Fig. 2 b]. Such thermal stability is in contrast with the light-induced H-CDW phase that was found to be suppressed above $T\!=\!60$ K Stojchevska et al. (2014). However, the C-CDW state can be recovered from the mixed state by warming up and slow-cooling across $T_{\mathrm{I-NC}}$ (Fig. 2 c). Subsequent thermal quenches from $T\!=\!300$ K do not recover the mixed state (Fig. 2 d), which can only be recovered upon heating and quenching from above $T_{\mathrm{I-NC}}$ (Fig. 2 e). Thus, our data confirms the IC-NC CDW as the necessary energy scale to access the mixed-CDW phase containing both the C-CDW and H-CDW order parameters.
Figure 3: Electronic structure of the mixed-CDW phase a -b Representative topography of the mixed-CDW phase showing domains, atomic resolution and the $\sqrt{13}\times\sqrt{13}$ reconstruction. c $dI/dV$ curves at each of the 6 domains seen in d. $dI/dV$ data reflects a narrower gap in the electronic structure of $1T$-TaS2 compared to the slow-cooled C-CDW phase (grey). Metallic behavior is observed in domain walls as indicated by point 3. d Resistivity, normalized to that at $T\!=\!300$ K for a mixed-sample (solid line) and as-grown sample (dotted grey). The inset shows a zoom in view around the TH-CDW/C-CDW transition.
The stability of the H-CDW phase at temperatures below $T_{\mathrm{I-NC}}$ in the mixed CDW phase enables us to ex-situ engineer the CDW mixed state at high temperatures (see Methods). We note that the mixed CDW state is stable over months, independent of the number of cleaves, and has been reproduced in quenched samples over multiple growth batches. Detailed temperature-dependent XRD measurements (See Methods) are shown in Fig. 2 f. All ex-situ quenched samples display twelve satellite reflections characteristic of the mixed-CDW state at low temperatures. Upon warming above $T^{WU}\!=\!210$ K, or cooling below $T^{CD}\!=\!195$ K, we observe a transition into an NC-CDW state with CDW Bragg peaks of two opposite chiralities. The existence of two chiral states in the NC-CDW is consistent with previous ultrafast electron diffraction experiments in $1T$-TaS2 driven with a low fluence optical pump reporting that the chirality of the low-temperature C-CDW domains is inherited from the symmetry-breaking IC-NC transition Haupt et al. (2016). We remark that this is distinct from the metastable H-CDW phase, which directly emerges from the C-CDW while maintaining the same chirality Stahl et al. (2020); Stojchevska et al. (2014); Gerasimenko et al. (2019). Detailed temperature dependence of the H-CDW-like peak and its NC-CDW precursor are shown in Fig. 2 h and i. C-CDW and H-CDW-like Bragg peaks show the same temperature dependence (see Extended data) suggesting the coupled nature of the two order parameters in the mixed-CDW phase. The asymmetrically narrower hysteretic loop of the first-order transition, when compared to as-grown samples, $T^{CD}\!=\!180$ K and $T^{WU}\!=\!210$ K, in conjunction with the narrower CDW Bragg peaks along L and the temperature dependence of the [0010] structural Brag peak shown in Fig. 2 c, suggests longer range interlayer ordering and a more 3D nature of the H-CDW phase.
Scanning tunneling microscopy (STM), measurements confirm the multi-domain nature of our quenched samples. Figure 3 a shows a characteristic $80\times 80$ nm2 area in a quenched sample. Each domain shows the atomic reconstruction of the sulfur layer due to low-temperature CDW, as shown in Fig. 3 c with $V\!=\!380$ mV, $I\!=\!0.35$ nA, scale bar length $\!=\!2$ nm). We note that the in-plane momentum resolution of our STM experiments prevents us from distinguishing between the C-CDW and the H-CDW-like domains. However, from the Fourier transform of STM images with atomic resolution, we extract a rotated satellite peak $\phi\!=\!12.8(5)^{\circ}$, that reflects the presence of domains of opposite chirality (See extended data).
To summarize, intermediate-quenched bulk $1T$-TaS2 samples host a mixed CDW state with two distinct order parameters, locked to opposite chiral phases, which emerge from a multi-chiral NC-CDW phase. Moreover, our temperature-dependent data reveals the narrowing of the hysteretic window across the NC-mixed-CDW phase transition. This indicates a stronger out-of-plane electronic correlation in the mixed-CDW phase of $1T$-TaS2.
Electronic structure of the bulk quenched samples. We now bring attention to the low-temperature electronic structure of the mixed-CDW phase. While the origin of the insulating nature of the C-CDW remains controversial Petocchi et al. (2022); Wang et al. (2020), it is well established that the c-axis stacking of CDW domains strongly influences the electronic structure of $1T$-TaS2 Ritschel et al. (2018); Lee et al. (2019); Butler et al. (2020); Ritschel et al. (2015). Moreover, recent STM and scanning tunneling spectroscopy (STS) studies of the stack shift between in-plane domains in consecutive layers have identified interlayer coupling as the key ingredient that determines the band structure of the low-temperature CDW phase in $1T$-TaS2 Cho et al. (2017); Ma et al. (2016); Park et al. (2021); Gerasimenko et al. (2019); Park et al. (2023). Thus, we use STS to study the low-temperature electronic state of the quenched samples. $dI/dV$ curves taken at $T\!=\!10$ K within six different domains, Fig. 3 a, reveal a reduction of the $E_{g}\!=\!300$ meV gap of the C-CDW Campbell et al. (2024) towards a state with a reduced but non-zero density of states at the Fermi level (Fig. 3 c). While a transition to a semimetallic state is consistent with that observed for the H-CDW Ma et al. (2016); Gerasimenko et al. (2019), the density of states maintains the two pronounced peaks at around $V_{B}=-300$ meV and $V_{B}=150$ meV, that have been associated with the lower and upper Hubbard bands Perfetti et al. (2005); Kim et al. (1994); Cho et al. (2015). Despite the finite in-plane density of states at $E_{F}$ and the formation of metallic domain walls (Fig. 3 b, curve 3), electrical transport in the mixed-CDW sample resembles that of the as-grown crystals.
We observed that quenched bulk samples display insulating behavior at all temperatures with the same relative resistivity increase with respect to $R(300$ K) as the C-CDW phase Thompson et al. (1971) (Figure 3 d). The mixed state resistivity exhibits a narrow thermal hysteresis that tracks that of the structural transition (Fig. 2 g), reflecting the role of out-of-plane order in the resistivity of $1T$-TaS2. The insulating nature of the mixed-CDW phase is distinct from the substantial $T<70$ K resistivity drop observed for the light-induced H-CDW phase in thin samples Stojchevska et al. (2014). This difference likely originates from the insulating C-CDW domains that occupy a large volume fraction of the mixed-CDW state and stacking CDW disorder between the two CDW orders with different periodicity along L (Figure 2 g-h). Additionally, we find that in-situ thermal cycling above $T_{IC}\!=\!350$  K and slow cooling recovers the characteristic resistivity of bulk $1T$-TaS2 with a hysteretical transition separated by $\Delta T\!=\!50$ K (Figure 3 d), consistent with our diffraction measurements in as-grown samples. Thus, our STS, resistivity, and XRD data support an interpretation in which bulk metallic transport is impeded by the stacking disorder along L of two order parameters with different long-range intralayer CDW order Liu et al. (2024).
Figure 4: Dynamic phase transition with two competing order parameters in $1T$-TaS2. a Schematic free energy landscape for a dynamical phase transition with two competing order parameters. A quench can stabilize a local minimum, different than the global minimum with order parameter $\psi_{1}$, corresponding to a secondary order parameter $\psi_{2}$. b-d Schematic of the possible low-temperature groundstates of $1T$-TaS2 as a function of quenching rate. c C-CDW, d mixed-CDW state and e H-CDW.
III Discussion
In summary, our results demonstrate that a mixed-CDW phase characterized by C-CDW and H-CDW domains of opposite chirality can be stabilized by a thermal quench through $T_{\mathrm{I-NC}}$. Our STS data reveals a semimetallic in-plane electronic density of states, while out-of-plane CDW stacking disorder prevents bulk metallic behavior.
It is compelling to compare our observations with theories for first-order dynamical phase transitions with two competing order parameters Sun and Millis (2020), in which the order parameter of the C-CDW corresponds to the global minimum of the free energy landscape. Under this hypothesis, the H-CDW phase is a local minimum of the free-energy landscape as sketched in Fig. 4 a. This provides a natural explanation for the large C-CDW domains for slow quenches, as sketched in Fig. 4 b and c. Such a free energy landscape allows a mixed inhomogeneous phase to be accessed at intermediate but adiabatic quenching rates in which a single temperature can be defined for all the different degrees of freedom in the material. This state is characterized by the coexistence of trapped regions of the two competing orders Del Re et al. (2016) and enhanced lifetime due to the transiently created domain walls Sun and Millis (2020). Our data is consistent with such a scenario: in $1T$-TaS2 a thermal quench stabilizes a mixed phase of H-CDW and C-CDW order parameters up to $T\!=\!180$ K in cooling and $T\!=\!210$ K in warming (Fig. 4 d).
We note that all our thermal quenching attempts were unsuccessful in engineering a sample without signatures of the C-CDW state, as sketched in Fig. 4 e. It is likely that the quench protocol we describe here results in a nonuniform quench rate that is locally too slow over some regions of the sample to allow for the nucleation of the H-CDw domains. Within this context, previous ultrafast stabilization of the H-phase in thin flakes Stojchevska et al. (2014); Stahl et al. (2020); Vaskivskyi et al. (2015); Hollander et al. (2015); Ma et al. (2016); Cho et al. (2016); Venturini et al. (2022); Tsen et al. (2015), can be understood as a non-thermal ultrafast quench Sun and Millis (2020). Our results are consistent with proposals in which the absorption of energy by the electronic degrees of freedom suppresses the C-CDW order parameter. The consequent relaxation of the non-thermal electrons via the lattice leads to an ionic displacement from that of the C-CDW state into that of the H-CDW Maklar et al. (2023) allowing for the correlations of the H-CDW order parameter to grow Sun and Millis (2020). Thus, our results suggest that the minimum of the free energy landscape associated with the H-CDW order parameter already exists at the thermodynamic equilibrium of $1T$-TaS2 and can be accessed with a sufficiently fast adiabatic quench (Figure 3 a).
Our quenching protocol and the observation of the mixed-CDW phase opens a new pathway towards the development of new efficient functional devices based on $1T$-TaS2 in which H and C-CDW domain engineering can lead to switchable metallic behavior at temperatures above LN2 Mihailovic et al. (2021); Mraz et al. (2022); Venturini et al. (2022). More generally, our results have implications for the dynamic control of quantum materials de la Torre et al. (2021); Jarc et al. (2023) by establishing that for macroscopic order parameters, the engineering of free energy landscapes can be achieved without the need for strong laser pulses when a thermally stable phase is already present at equilibrium Chandran et al. (2012); Du et al. (2023); Wu et al. (2024).
IV Data Availability
The datasets obtained in this study are available in the open database at (To be provided).
V Acknowledgements
A.d.l.T acknowledges helpful conversation with Michael Buchold, Yue Cao, Martin Claassen, Greg Fiete, Simon Gerber, and Dante M. Kennes. Work performed at Brown University by A.d.l.T., Q.W. and K.W.P. was supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, under Award Number DE-SC0021223. P.M.V. and D.B. acknowledge support from NSF DMR 2226097 and the Mason Graduate Division’s Presidential Scholarship Program”. This work is based on research conducted at the Center for High-Energy X-ray Sciences (CHEXS), which is supported by the National Science Foundation (BIO, ENG and MPS Directorates) under award DMR-1829070.
Author Contributions
A.d.l.T and Q. W. synthesized and characterized the single crystal samples. A.d.l.T., Q. W., K. W. P. and J.P.C.R. performed the High Dynamic Range x-ray scattering measurements. All x-ray experimental data was analyzed and interpreted by A.d.l.T and K .W. P.. B. C., J. V. R., and S. M. H. performed and analyzed the STM and STS measurements. D. B. and P.M.V. performed the cryo-Raman experiments.‘ A.d.l.T. and. K. W. P wrote the manuscript with input from all co-authors.
VI Methods
VI.1 Sample Growth
$1T$-TaS2 bulk single crystals were grown by chemical vapor transport, with iodine as the transport agent. Stoichiometric amounts of elemental Ta and S were sealed in a quartz tube. The tubes were held at a $950^{\circ}$ C – $850^{\circ}$ C temperature gradient for 240 h and then placed in ice water to stabilize the 1T phase. Samples were characterized by XRD, resistivity, and Raman (see Fig. 2).
VI.2 Thermal quench
Bulk $1T$-TaS2 crystals were sealed in a 1" quartz tube filled with Ar gas. The $1T$-TaS2 samples were then annealed by warming them up to 420 K, below the polytype transition at 660 K, with an 8h hold. This was followed by either a fast quench in which the quartz tube is dropped in a reservoir of LN2 or slow cooling at room temperature. In-situ quench was performed at the QM2 beamline of the CHESS synchrotron by alternating between hot gas and a LN2 cryo-stream.
VI.3 Estimate of the cooling rate
To provide an estimate of the cooling rate for our $1T$-TaS2 sample quenched within an Ar-filled quartz ampule shown in Fig. 4 a and Extended data, we solve the heat diffusion equation for a system with cylindrical symmetry under the assumption that $1T$-TaS2 is at all times thermalized with the Ar bath and that the outside surface of the quartz tube is thermalized with the bath. Thus, the limiting factor to thermalization is the heat diffusion from the hot Ar volume at the starting temperature ($\rho{Ar}=1.8$ kg/m3, $K_{Ar}=0.0179$ W/(m·K), $C_{p}^{Ar}=400$ Jkg-1m) through the $1.1$ mm thick quartz wall with the outer surface acting as the heat sink ($\rho_{Q}=2650$ kg/$m^{3}$, $K_{Q}=1$,$C_{p}^{Q}=700$ JKg-1m)).
VI.4 X-ray High Dynamic Range Reciprocal Space Mapping
Hard x-ray (20 keV) HDRM was performed at the QM2 beamline of Cornell High Energy Synchrotron Source (CHESS) using a Pilatus 6M area detector. The sample is mounted on a Kapton pin using GE varnish and rotated around three different axes by 360. Data is collected at every 0.1 step with a frame rate of 0.1s. The data is reduced into a 3D stack, indexed, and projected into a 2D dataset using the beamline software.
VI.5 In-house XRD measurements
In-house XRD measurements were performed on a Huber four-circle diffractometer in combination with a micro-focused Ag $K_{\alpha}$ x-ray source and a 256 x 256 pixel GaAs detector (pixel size 55 x 55 $\mu$m) situated 96 cm away from the center of rotation. Data was reduced using the same software as for the CHESS data.
VI.6 STM/STS measurements
STM and STS measurements were performed on a RHK Technology PanScan Freedom closed-cycle STM. Samples for the STM/STS measurements were mounted on stainless steel post holders using conductive silver epoxy and quenched as previously described. After the quench, the samples were characterized by XRD and kept below $T=370$K. The surface of a $1T$-TaS2 crystal was cleaved at room temperature and $\approx 10^{-10}$ mbar by carbon tape and a cleaving screw before being introduced to the STM stage and cooled to 10 K. All data presented here were taken at 10 K.
VI.7 Resistivity
Resistivity measurements were performed using a Quantum Design Physical Properties Measurement System (PPMS). Samples were mounted on a DC resistivity puck and a standard four-probe method was used, with Au wires and Ag paste contacts.
VI.8 Raman Spectroscopy
Raman measurements of quenched and unquenched TaS2 single crystals are performed at 5 K in a closed-cycle He cryostat. The samples are excited in a backscattering geometry with a 532 nm laser focused through a 0.6 NA objective lens with 40$\times$ magnification. The laser power was measured to be 110 $\mu$W before the objective. The scattered light was collected through the same lens and directed to an imaging spectrograph with a liquid nitrogen-cooled charge-coupled device. Rejection of the Rayleigh scattered light is accomplished using three-volume Bragg gratings that allow measurement of Raman photons down to 10 cm-1.

### v3

@@
 Dynamic phase transition into a mixed-CDW state in 1T-TaS2 via a thermal quench
-A.
-de la Torre a.delatorreduran@northeastern.edu Quantum Materials and Sensing Institute, Northeastern University, Burlington, Massachusetts, 01803, USA Department of Physics, Northeastern University, Boston, Massachusetts, 02115, USA Department of Physics, Brown University, Providence, Rhode Island 02912, United States    Q.
-Wang Department of Physics, Brown University, Providence, Rhode Island 02912, United States    B.
-Campbell Department of Physics and Astronomy, University of New Hampshire, Durham, NH 03824.
-J.
-V.
+Alberto de la Torre a.delatorreduran@northeastern.edu Department of Physics, Northeastern University, Boston, Massachusetts, 02115, USA Quantum Materials and Sensing Institute, Northeastern University, Burlington, Massachusetts, 01803, USA Department of Physics, Brown University, Providence, Rhode Island 02912, United States    Qiaochu Wang Department of Physics, Brown University, Providence, Rhode Island 02912, United States    Yasamin Masoumi Department of Physics, Northeastern University, Boston, Massachusetts, 02115, USA    Benjamin Campbell Department of Physics and Astronomy, University of New Hampshire, Durham, NH 03824.
+Jake V.
 Riffle Department of Physics and Astronomy, University of New Hampshire, Durham, NH 03824.
-D.
-Balasundaram Department of Physics and Astronomy, George Mason University, Fairfax, VA, 22030 USA Quantum Science and Engineering Center, George Mason University, Fairfax, VA, 22030 USA    P.M.Vora Department of Physics and Astronomy, George Mason University, Fairfax, VA, 22030 USA Quantum Science and Engineering Center, George Mason University, Fairfax, VA, 22030 USA    J.P.C.
-Ruff Cornell High Energy Synchrotron Source, Cornell University, Ithaca, NY, 14853, USA    S.
-M.
+Dushyanthini Balasundaram Department of Physics and Astronomy, George Mason University, Fairfax, VA, 22030 USA Quantum Science and Engineering Center, George Mason University, Fairfax, VA, 22030 USA    Pattrick M.Vora Department of Physics and Astronomy, George Mason University, Fairfax, VA, 22030 USA Quantum Science and Engineering Center, George Mason University, Fairfax, VA, 22030 USA    Jacob P.C.
+Ruff Cornell High Energy Synchrotron Source, Cornell University, Ithaca, NY, 14853, USA    Gregory A.
+Fiete Department of Physics, Northeastern University, Boston, Massachusetts, 02115, USA Quantum Materials and Sensing Institute, Northeastern University, Burlington, Massachusetts, 01803, USA Department of Physics, Massachusetts Institute of Technology, Cambridge, MA 02139, USA    Shawna M.
 Hollen Department of Physics and Astronomy, University of New Hampshire, Durham, NH 03824.
-K.
-W.
+Kemp W.
 Plumb kemp˙plumb@brown.edu Department of Physics, Brown University, Providence, Rhode Island 02912, United States
-(July 12, 2024)
+(January 8, 2025)
 Ultrafast light-matter interaction has emerged as a new mechanism to exert control over the macroscopic properties of quantum materials toward novel functionality.
-To date, technological applications of these non-thermal phases are limited by their ultrashort lifetimes and low-ordering temperatures.
-Among the most studied photoinduced metastable phases for their technological promise is the hidden metallic charge density wave (H-CDW) in the model correlated CDW compound $1T$-TaS2.
-Despite active study and engineering, the nature of the photoinduced H-CDW remains the subject of debate and is only accessible at cryogenic temperatures.
-Here, we stabilize the H-CDW phase at thermal equilibrium up to near-room temperature by accessing an intermediate mixed CDW order regime via thermal quenching.
-Using x-ray high dynamic range reciprocal space mapping (HDRM) and scanning tunneling spectroscopy (STS), we reveal the coexistence of commensurate (C) CDW and H-CDW domains below 180 K during cooling and below 210 K during warming.
+To date, technological applications of non-thermal phases are limited by their ultrashort lifetimes and low-ordering temperatures.
+The hidden metallic charge density wave (H-CDW) in $1T$-TaS2 is among the most studied photoinduced metastable phases because of its technological promise.
+However, despite active study, the nature of the photoinduced H-CDW remains the subject of debate and potential applications have been limited because it has so far only been stabilized at cryogenic temperatures.
+Here, we stabilize the H-CDW phase at thermal equilibrium by accessing a mixed CDW order regime via thermal quenching.
+Using x-ray high dynamic range reciprocal space mapping (HDRM) and scanning tunneling spectroscopy (STS), we reveal the coexistence of commensurate (C) CDW and H-CDW domains up to 210 K.
 Our findings show that each order parameter breaks basal plane mirror symmetry with different chiral orientations and induces out-of-plane unit cell tripling in the H-CDW phase.
 Despite metallic domain walls and a finite density of states at zero bias observed via STS, bulk resistance remains insulating due to CDW stacking disorder.
-This study establishes the H-CDW as a thermally stable phase and introduces a new mechanism for switchable metallic behavior in thin flakes of $1T$-TaS2 and similar materials with competing order phases.
+By comparing our data to Landau-Ginzburg theory calculations of the free energy functional, our study establishes the H-CDW as a thermally stable phase and introduces a new mechanism for switchable metallic behavior in thin flakes of $1T$-TaS2 and similar materials with competing phases.
 I Introduction
-The correlated van der Waal material $1T$-TaS2 displays a series of first-order phase transitions associated with the onset of CDW orders with different degrees of commensurability Thompson et al.
+The correlated van der Waals material $1T$-TaS2 displays a series of first-order phase transitions associated with the onset of CDW orders with different degrees of commensurability Thompson et al.
 (1971); Wilson et al.
 (1975); Fazekas and Tosatti (1979).
-First, at $T=550$ K, a mirror-symmetry preserving incommensurate (I) CDW onsets from a metallic high-temperature state Sipos et al.
+Lowering from high temperatures, at $T=550$ K a mirror-symmetry preserving incommensurate (I) CDW onsets from a metallic state Sipos et al.
 (2008).
 At $T_{\mathrm{I-NC}}$$\approx 350$ K, $1T$-TaS2 enters into a nearly commensurate (NC) CDW phase.
-The atomic rearrangement associated with the incommensurate NC-CDW (in-plane ordering wavevector $q_{||}\!=\!0.284$ r.l.u.) leads to disconnected regions displaying a $\sqrt{13}\times\sqrt{13}$ structural reconstruction in which twelve Ta atoms move towards a central one forming a star-of-David (SOD) pattern Spijkerman et al.
+The atomic rearrangement associated with the incommensurate NC-CDW (in-plane ordering wavevector $q_{||}\!=\!0.284$ r.l.u.) leads to disconnected regions displaying a $\sqrt{13}\times\sqrt{13}$ spatial reconstruction in which twelve Ta atoms move towards a central one forming a star-of-David (SOD) pattern Spijkerman et al.
 (1997).
-The NC-CDW breaks the mirror plane symmetries of the lattice basal plane and spontaneously selects one of two chiral domains rotated from the $a$-axis by $\phi\approx 12^{\circ}$.
+The NC-CDW breaks the mirror symmetries of the lattice basal plane and spontaneously selects one of two chiral domains rotated from the $a$-axis by $\phi\approx 12^{\circ}$.
 Then, at $T_{\mathrm{NC-C}}\approx 175$ K in cooling and $T_{\mathrm{NC-C}}\approx 225$ K in warming, a 3D CDW order emerges from the NC-CDW phase.
@@
 (1975).
-In general, experimental probes spanning a large range of lengthscales find that $1T$-TaS2 samples lack domain twinning that is expected across a first-order phase transition Zong et al.
+Experimental probes spanning a large range of length scales find that in the C-CDW phase of $1T$-TaS2 samples lack the chiral twin domains expected across the $T_{\mathrm{NC-C}}$ first-order phase transition Zong et al.
 (2018); Fichera et al.
@@
 (2024).
-This reflects the kinetics of the first cooling across the only first-order phase transition in $1T$-TaS2 that breaks in-plane mirror symmetries, $T_{\mathrm{I-NC}}$ Sept and Tuszyński (1994); Binder (1987).
-During synthesis, the $1T$ phase is stabilized by quenching to $T\approx 273$ K across the polytype transition, $T_{\rm{1T-2H}}\!=\!600$ K Thompson et al.
+The stabilization of a single chiral domain reflects the kinetics of the cooling across the only first-order phase transition in $1T$-TaS2 that breaks in-plane mirror symmetries, $T_{\mathrm{I-NC}}$ Sept and Tuszyński (1994); Binder (1987).
+During synthesis, the metastable $1T$ phase is stabilized by quenching from $T\approx 1000$ K to $T\approx 273$ K across the polytype transition, $T_{\rm{1T-2H}}\!=\!600$ K Thompson et al.
 (1971).
-During this initial quench, the cooling rate across $T_{\mathrm{I-NC}}$ is effectively slow when compared to the intrinsic timescales of the electronic ($\approx$ fs) and atomic ($\approx$ ps) degrees of freedom and enables the nucleation of large chiral CDW domains with lateral sizes comparable to the correlation length of the order parameter parameter Oxtoby (1998).
-The chiral domain structure of the C-CDW in $1T$-TaS2 is stable to cooling through $T_{\mathrm{NC-C}}$ Zong et al.
+During this initial quench, the cooling rate across $T_{\mathrm{I-NC}}$ is effectively slow when compared to the intrinsic timescales of the electronic ($\approx$ fs) and lattice ($\approx$ ps) degrees of freedom and enables the nucleation of large chiral CDW domains with lateral sizes comparable to the correlation length of the order parameter parameter Oxtoby (1998).
+The chiral mono-domain structure of the C-CDW in $1T$-TaS2 is stable to cooling through $T_{\mathrm{NC-C}}$ Zong et al.
 (2018); Fichera et al.
@@
 (2023); Zong et al.
-(2018).Thus, $T_{\mathrm{I-NC}}$ emerges as the relevant energy scale for chiral CDW domain nucleation in $1T$-TaS2 as the initial mirror-symmetry breaking transition in the CDW phase diagram Thompson et al.
 (2018).
-Moreover, x-ray pair distribution function has revealed that the in-plane Ta average displacement from the mean crystallographic position, a measure of the CDW formation, remains constant across the NC-C CDW transition.
-At the same time, it dramatically changes across $T_{\mathrm{I-NC}}$ Bozin et al.
+Thus, $T_{\mathrm{I-NC}}$ emerges as the relevant energy scale for chiral CDW domain nucleation in $1T$-TaS2 as the temperature at which the initial mirror-symmetry breaking transition in the CDW phase diagram Thompson et al.
+(1971) occurs and the largest specific heat divergence is observed Mañas-Valero et al.
+(2021).
+Moreover, x-ray pair distribution function measurements have revealed that the anisotropic Ta Debye-Waller factor, associated with CDW formations, remains constants across the NC-C CDW transition but dramatically changes across $T_{\mathrm{I-NC}}$ Bozin et al.
 (2023).
@@
 (2016); Maklar et al.
-(2023) These domains maintain the chirality of the C-CDW Stahl et al.
+(2023).
+These domains maintain the chirality of the C-CDW Stahl et al.
 (2020) and are related by a relative translation of the SOD pattern by a finite number of atomic lattice vectors Ma et al.
@@
 (2021).
-Moreover, slight differences between the displacement of the Ta atoms between the H-CDW and the C-CDW distinguish the two.
+Slight differences between the displacement of the Ta atoms between the H-CDW and the C-CDW distinguish the two.
 The H-CDW phases result in CDW Bragg peaks rotated by $\phi\!=\!11.9^{\circ}$ from the $a^{*}$-axis with a larger basal plane ordering wavevector, $q_{||}\!=\!0.284$ r.l.u..
@@
 (2024).
-However, the H-CDW has only been stabilized for ultrashort periods (${}100\mu$s) at temperatures above $T\!=\!50$ K Stojchevska et al.
+However, the H-CDW has only been stabilized for ultrashort periods ($\sim 100\,\,\mu$s) at temperatures above $T\!=\!50$ K Stojchevska et al.
 (2014); Ravnik et al.
@@
 (2017).
-Here, we demonstrate that the H-CDW phase exists at thermal equilibrium in $1T$-TaS2 and extend its thermal stability to near-room temperature by accessing a new long-lived mixed-CDW state via a reversible thermal quench across $T_{\mathrm{I-NC}}$.
-We deploy x-ray high dynamic range reciprocal space mapping (HDRM) to unequivocally demonstrate the co-existence of the C-CDW and H-CDW states in thermally quenched samples.
+Here, we demonstrate that the H-CDW phase exists in thermal equilibrium in $1T$-TaS2 and extend its thermal stability by a factor of three by accessing a new long-lived mixed-CDW state via a reversible thermal quench across $T_{\mathrm{I-NC}}$.
+We use x-ray high dynamic range reciprocal space mapping (HDRM) to unequivocally demonstrate the co-existence of the C-CDW and H-CDW states in thermally quenched samples.
 Each CDW order parameter is linked to a mirror-symmetry breaking chiral domain.
 The mixed-CDW state emerges from a multi-domain NC-CDW state below $T^{CD}=180$ K in cooling and $T^{WU}=210$ K in warming.
-The mixed-CDW remains stable up to $T_{\mathrm{I-NC}}$.
+The mixed-CDW low temperature state can be recovered as long as the system is not heated above $T_{I-NC}$.
 Scanning tunneling spectroscopy (STS) shows that the mixed-CDW is characterized by a semi-metallic in-plane density of states.
 At the same time, bulk resistivity remains insulating due to out-of-plane CDW stacking disorder in mixed CDW bulk samples.
-Our results demonstrate a new state in the phase diagram of $1T$-TaS2, enabling a new mechanism to engineer a semimetallic density of states at near room temperature.
+Our results and calculations of the free energy landscape within the Ginzburg-Landau theory demonstrate a new thermally accessible state in the phase diagram of $1T$-TaS2, enabling a new mechanism to engineer a semimetallic density of states at temperatures well above liquid nitrogen temperatures.
 […]
@@
 Identical reciprocal space maps were obtained across many cooling cycles and samples.
-However, we found that the C-CDW domain structure depends on the cooling rate, and a fast thermal quench across $T_{\mathrm{I-NC}}$ can consistently produce a multi-domain CDW.
+We found that the C-CDW domain structure in a given sample depends on the cooling rate through $T_{\mathrm{I-NC}}$ and fast thermal quench across $T_{\mathrm{I-NC}}$ stabilizes a multi-chiral domain CDW.
 Fig.
 1 b shows the [HK0] reciprocal space map after an in-situ quench in our bulk sample across $T_{\mathrm{I-NC}}$ down to $T\!=\!140$ K at a rate of $0.2$ K/ms near $T_{\mathrm{I-NC}}$ (See Methods for a description of the in-situ quench method and calculation of quenching rates using heat diffusion calculations).
-We found that thermal quenching consistently resulted in a new set of superlattice reflections of opposite chirality to the original set of satellite peaks ($\phi\!=\!-12.3^{\circ}$) [Fig.
-1 b].
+We found that thermal quenching consistently resulted in a new set of superlattice reflections of opposite chirality to the original set of satellite peaks ($\phi\!=\!-12.3^{\circ}$) (Fig.
+1 b, highlighted by blue markers).
 The superior reciprocal space resolution of synchrotron single crystal HDRM enables us to resolve differences in the wavevector of the two sets of CDW Bragg peaks with different chirality.
-The in-plane wavevector magnitude $|q_{||}|\!=\!0.284(3)~{}\AA^{-1}$ of the quenched-induced satellite peaks is larger than that of the C-CDW ($|q_{||}|\!=\!0.277(3)~{}\AA^{-1}$) and agrees with that of the light-stabilized H-CDW state Stahl et al.
+The in-plane wavevector magnitude $|q_{||}|\!=\!0.284(3)~{}\AA^{-1}$ of the quenched-induced satellite peaks is larger than that of the C-CDW ($|q_{||}|\!=\!0.277(3)~{}\AA^{-1}$) and is the same as the light-stabilized H-CDW state Stahl et al.
 (2020).
@@
 We also find that the c-axis periodicity differs between the C-CDW and H-CDW-like domains.
-Figure 1 d and e show [H0L] reciprocal space maps centered at H = 1 r.l.u and L = 0 r.l.u before and after the thermal quench, respectively.
+Figure 1 d and e show [H0L] reciprocal space maps centered at H = 1 r.l.u.
+and L = 0 r.l.u.
+($\delta k=0.3$ r.l.u.) before and after the thermal quench, respectively.
 The as-grown sample displays a broad two-peak structure, with the center of mass at integer values of L.
-These short-range out-of-plane correlations are characteristic of the dimerization of CDW layers in the low-temperature C-CDW phase (Fig.
-1 g) Tanda et al.
+These short-range out-of-plane correlations are characteristic of the dimerization of CDW layers in the low-temperature C-CDW phase Tanda et al.
 (1984).
-On the other hand, the H-CDW-like satellite peaks shift towards $l\!=\!\pm 1/3$ indicating a collapse of the interlayer dimerization towards a triple layer CDW stacking (Fig.
-1 h) as was observed for the light-induced H-CDW Stahl et al.
+On the other hand, while retaining a dumbbell-like shape, the H-CDW-like satellite peaks shift towards $l\!=\!\pm 1/3$ indicating a collapse of the interlayer dimerization towards a triple layer CDW stacking as was observed for the light-induced H-CDW Stahl et al.
 (2020).
@@
 1 f).
-Our HDRM data in quenched samples demonstrates the coexistence of the C-CDW with a second CDW phase with the same order parameter as the H-CDW phase at much higher temperatures ($T=140$ K).
+Similar sharpening of the out of plane CDW peaks was also observed in light-induced H-CDW $1T$-TaS2 Stahl et al.
+(2020).
+Altogether, our HDRM data in quenched samples demonstrates the coexistence of the C-CDW with a second CDW phase with the same order parameter as the H-CDW phase at much higher temperatures ($T=140$ K).
 Critical temperature and metastability of the quenched mixed-CDW state.
@@
 The inset shows a zoom in view around the TH-CDW/C-CDW transition.
-The stability of the H-CDW phase at temperatures below $T_{\mathrm{I-NC}}$ in the mixed CDW phase enables us to ex-situ engineer the CDW mixed state at high temperatures (see Methods).
+The possibility of recovering the H-CDW phase as part of the stable mixed-CDW phase by keeping the samples below $T_{\mathrm{I-NC}}$ enables us to ex-situ form the CDW mixed state (see Methods) and perform additional characterization of this state.
 We note that the mixed CDW state is stable over months, independent of the number of cleaves, and has been reproduced in quenched samples over multiple growth batches.
-Detailed temperature-dependent XRD measurements (See Methods) are shown in Fig.
-2 f.
-All ex-situ quenched samples display twelve satellite reflections characteristic of the mixed-CDW state at low temperatures.
-Upon warming above $T^{WU}\!=\!210$ K, or cooling below $T^{CD}\!=\!195$ K, we observe a transition into an NC-CDW state with CDW Bragg peaks of two opposite chiralities.
-The existence of two chiral states in the NC-CDW is consistent with previous ultrafast electron diffraction experiments in $1T$-TaS2 driven with a low fluence optical pump reporting that the chirality of the low-temperature C-CDW domains is inherited from the symmetry-breaking IC-NC transition Haupt et al.
+Detailed temperature-dependent XRD measurements (See Methods) of ex-situ quenched samples display twelve satellite reflections characteristic of the mixed-CDW state at low temperatures (Fig.2 f).
+For temperatures above $T^{WU}_{NC-Mixed}\!=\!210$ K during warming up, or below $T^{CD}_{NC-mixed}\!=\!195$ K during cooling down, we observe a transition into the NC-CDW.
+However, the NC-CDW state in quenched samples is characterized by two sets of CDW Bragg peaks with opposite chiralities.
+This state is shown in Fig.
+2 f at $T=270$ K for ex-situ quenched samples.
+Equivalent HDRM data at $T=300$ K is shown in Extended Data 2.
+We note that the L dependence of the two NC-CDW satellite peaks is different from that of the H-CDW we observe at low temperature (see Extended data 2), ruling out a supercooled NC-CDW phase in our quenched samples Yoshida et al.
 (2015).
-We remark that this is distinct from the metastable H-CDW phase, which directly emerges from the C-CDW while maintaining the same chirality Stahl et al.
+The existence of two chiral domains in the NC-CDW is consistent with previous ultrafast electron diffraction experiments that demonstrated the chirality of the low-temperature C-CDW domains was inherited from the symmetry-breaking at the IC-NC transition Haupt et al.
+(2016) and that two NC-CDW domains were stabilized after laser excitation at room temperature Zong et al.
+(2018).
+A detailed temperature dependence of the H-CDW-like and NC-CDW Bragg reflections are shown in Fig.
+2 h and i.
+Upon stabilizing the mixed state at low temperatures via a quench across $T_{I-NC}$, the H-CDW peaks directly emerge from NC-CDW peaks of corresponding chirality.
+This temperature dependence is distinct from the metastable H-CDW phase, which is suppressed above $T=60$ K.
+Stahl et al.
 (2020); Stojchevska et al.
@@
 (2019).
-Detailed temperature dependence of the H-CDW-like peak and its NC-CDW precursor are shown in Fig.
-2 h and i.
-C-CDW and H-CDW-like Bragg peaks show the same temperature dependence (see Extended data) suggesting the coupled nature of the two order parameters in the mixed-CDW phase.
-The asymmetrically narrower hysteretic loop of the first-order transition, when compared to as-grown samples, $T^{CD}\!=\!180$ K and $T^{WU}\!=\!210$ K, in conjunction with the narrower CDW Bragg peaks along L and the temperature dependence of the [0010] structural Brag peak shown in Fig.
-2 c, suggests longer range interlayer ordering and a more 3D nature of the H-CDW phase.
+We found that C-CDW and H-CDW-like Bragg peaks show the same temperature dependence (see Extended data) suggesting the coupled nature of the two order parameters in the mixed-CDW phase.
+The asymmetrically narrower hysteretic loop of the first-order transition, when compared to as-grown samples, $T^{CD}_{NC-C}\!=\!180$ K and $T^{WU}_{NC-C}\!=\!210$ K, in conjunction with the narrower CDW Bragg peaks along L and the temperature dependence of the [0010] structural Bragg peak shown in Fig.
+2 c, suggests longer range interlayer correlations and a more 3D nature of the H-CDW phase.
 Scanning tunneling microscopy (STM), measurements confirm the multi-domain nature of our quenched samples.
 Figure 3 a shows a characteristic $80\times 80$ nm2 area in a quenched sample.
-Each domain shows the atomic reconstruction of the sulfur layer due to low-temperature CDW, as shown in Fig.
+Each domain shows the atomic reconstruction of the sulfur layer due to a low-temperature CDW, as shown in Fig.
 3 c with $V\!=\!380$ mV, $I\!=\!0.35$ nA, scale bar length $\!=\!2$ nm).
 We note that the in-plane momentum resolution of our STM experiments prevents us from distinguishing between the C-CDW and the H-CDW-like domains.
-However, from the Fourier transform of STM images with atomic resolution, we extract a rotated satellite peak $\phi\!=\!12.8(5)^{\circ}$, that reflects the presence of domains of opposite chirality (See extended data).
+However, from the Fourier transform of STM images with atomic resolution, we extract a rotated satellite peak $\phi\!=\!12.8(5)^{\circ}$, that reflects the presence of domains of opposite chirality (related by a mirror operation along one of the high symmetry crystallographic directions) (See extended data).
+Our STM data resembles the in-plane inhomogeneity observed in optically and electrically driven $1T$-TaS2, where multiple domains of opposite phase have been reported Gerasimenko et al.
+(2019); Ma et al.
+(2016); Maklar et al.
+(2023).
+We note that the smaller domain density prevented us from performing a comprehensive in-plane phase configuration analysis of the chiral domains, but we cannot exclude the presence of phase domains (related by a translation operation) Gerasimenko et al.
+(2019); Ma et al.
+(2016); Maklar et al.
+(2023).
 To summarize, intermediate-quenched bulk $1T$-TaS2 samples host a mixed CDW state with two distinct order parameters, locked to opposite chiral phases, which emerge from a multi-chiral NC-CDW phase.
@@
 (2015).
-Moreover, recent STM and scanning tunneling spectroscopy (STS) studies of the stack shift between in-plane domains in consecutive layers have identified interlayer coupling as the key ingredient that determines the band structure of the low-temperature CDW phase in $1T$-TaS2 Cho et al.
+Moreover, recent STM and scanning tunneling spectroscopy (STS) studies of the stacking shift between in-plane domains in consecutive layers have identified interlayer coupling as the key ingredient that determines the band structure of the low-temperature CDW phase in $1T$-TaS2 Cho et al.
 (2017); Ma et al.
@@
 (2016); Gerasimenko et al.
-(2019), the density of states maintains the two pronounced peaks at around $V_{B}=-300$ meV and $V_{B}=150$ meV, that have been associated with the lower and upper Hubbard bands Perfetti et al.
+(2019), the density of states maintains two pronounced peaks at around $V_{B}=-300$ meV and $V_{B}=150$ meV, that have been associated with the lower and upper Hubbard bands Perfetti et al.
 (2005); Kim et al.
@@
 3 b, curve 3), electrical transport in the mixed-CDW sample resembles that of the as-grown crystals.
+Figure 4: Dynamic phase transition with two competing order parameters in $1T$-TaS2.
+a Cooling rate estimates for an optically induced quench (orange) and solutions of the heat diffusion equation for the slow (red) and fast (blue) cooling schemes used in this work (See Methods and Extended data).
+b Meanfield-only free energy functional for the two competing CDW orders, with a competition factor $d_{0}=0.8$, corresponding to an adiabatic quench.
+c 2D projection of the energy functional in b.
+d Meanfield-only free energy functional for the two competing CDW orders, with a competition factor $d_{0}=6.1$, corresponding to laser quench.
+Fast quenches correspond to larger values of $d_{0}$.
+e 2D projection of the energy functional in d.
+The yellow and red markers indicate local and global minima, respectively.
+f Computed phase diagram as a function of the competition factor, $d_{0}$, and mass term, $a_{0}$.
+The mixed CDW phase can only be stabilized in a narrow range of intermediate-rate quenches.
+Faster quenches (optical or electrical pulses) lead to the suppression of the C-CDW phase and the emergence of the H-CDW phase
 We observed that quenched bulk samples display insulating behavior at all temperatures with the same relative resistivity increase with respect to $R(300$ K) as the C-CDW phase Thompson et al.
@@
 (2024).
-Figure 4: Dynamic phase transition with two competing order parameters in $1T$-TaS2.
-a Schematic free energy landscape for a dynamical phase transition with two competing order parameters.
-A quench can stabilize a local minimum, different than the global minimum with order parameter $\psi_{1}$, corresponding to a secondary order parameter $\psi_{2}$.
-b-d Schematic of the possible low-temperature groundstates of $1T$-TaS2 as a function of quenching rate.
-c C-CDW, d mixed-CDW state and e H-CDW.
 III Discussion
 In summary, our results demonstrate that a mixed-CDW phase characterized by C-CDW and H-CDW domains of opposite chirality can be stabilized by a thermal quench through $T_{\mathrm{I-NC}}$.
-Our STS data reveals a semimetallic in-plane electronic density of states, while out-of-plane CDW stacking disorder prevents bulk metallic behavior.
-It is compelling to compare our observations with theories for first-order dynamical phase transitions with two competing order parameters Sun and Millis (2020), in which the order parameter of the C-CDW corresponds to the global minimum of the free energy landscape.
-Under this hypothesis, the H-CDW phase is a local minimum of the free-energy landscape as sketched in Fig.
-4 a.
-This provides a natural explanation for the large C-CDW domains for slow quenches, as sketched in Fig.
-4 b and c.
-Such a free energy landscape allows a mixed inhomogeneous phase to be accessed at intermediate but adiabatic quenching rates in which a single temperature can be defined for all the different degrees of freedom in the material.
+The required thermal quench is at an intermediate rate ($\approx 120$ K/s) based on heat diffusion calculations (see Methods), faster than the initial quench that $1T$-TaS2 crystals experience during growth ($\leq 2$ K/s) but slower than the changes to the electronic temperature induced by a laser pulse ($\geq 1$ K/fs) Zong et al.
+(2018); Stojchevska et al.
+(2014) (Figure 4) a.
+We note that the chiral splitting (relative rotation) of the two order parameters enables us to clearly resolve the mixed-CDW phase.
+If the H-CDW and C-CDW phases exhibited the same chirality, the small difference in their ordering wavevectors would lead to overlapping satellite peaks, complicating the observation of this novel state.
+Our STS data reveals that this mixed state hosts a semimetallic in-plane electronic density of states, while out-of-plane CDW stacking disorder prevents bulk metallic behavior.
+It is compelling to compare our observations with theories for dynamical phase transitions Sun and Millis (2020); Sefidkhani et al.
+(2024) and how domain formation influences the competition between coexisting order parameters.
+Thus, to describe the low temperature state after a quench, we consider a mean-field energy functional McMillan (1975) within the Landau-Ginzburg theory including, among others (see the Supplementary Information for more details of the calculations), a $d_{0}\bar{\phi}^{2}_{C}\bar{\phi}^{2}_{H}$ term in which $d_{0}$ parameterized the competition between the C-CDW and the H-CDW phase.
+The order parameter is the spatial charge modulation, $\alpha(\vec{r})=\Re(\phi_{C}e^{i\vec{Q}_{C}\cdot\vec{r}}+\phi_{H}e^{i\vec{Q}_{H}\cdot\vec{r}})$ in which the amplitude of the order parameter for the C-CDW, ${\phi}_{C}$, corresponds to the global minimum of the free energy landscape for slow cooling protocols, and $\phi_{H}$ is the amplitude of the H-CDW order parameter.
+We note that our model does not include the order parameter for the NC-CDW as the initial quench from $T_{\mathrm{I-NC}}$ down to low temperature effectively skips over the intermediate CDW state.
+Morevoer, the H-CDW phase can only be stabilized from the C-CDW phase Stojchevska et al.
+(2014) indicating that the NC-CDW is not necessary to discuss the competition between these two states.
+Such free energy landscape allows a mixed inhomogeneous phase to be accessed at intermediate but adiabatic quenching rates in which a single temperature can be defined for all the different degrees of freedom in the material.
 This state is characterized by the coexistence of trapped regions of the two competing orders Del Re et al.
 (2016) and enhanced lifetime due to the transiently created domain walls Sun and Millis (2020).
-Our data is consistent with such a scenario: in $1T$-TaS2 a thermal quench stabilizes a mixed phase of H-CDW and C-CDW order parameters up to $T\!=\!180$ K in cooling and $T\!=\!210$ K in warming (Fig.
-4 d).
-We note that all our thermal quenching attempts were unsuccessful in engineering a sample without signatures of the C-CDW state, as sketched in Fig.
-4 e.
+[…]
+Our data is consistent with a scenario in $1T$-TaS2 in which a thermal quench stabilizes a mixed phase of H-CDW and C-CDW order parameters up to $T\!=\!180$ K in cooling and $T\!=\!210$ K in warming.
+We note that coexistence of the H-CDW and C-CDW phase in ultrafast switched samples has also been recently reported Burri et al.
+(2024).
+Moreover, all our thermal quenching attempts were unsuccessful in engineering a sample without signatures of the C-CDW state indicating that sufficiently large values of $d_{0}$ where not achieved.
 It is likely that the quench protocol we describe here results in a nonuniform quench rate that is locally too slow over some regions of the sample to allow for the nucleation of the H-CDW domains.
@@
 (2022); Tsen et al.
-(2015), can be understood as a non-thermal ultrafast quench Sun and Millis (2020).
+(2015), can be understood as a non-thermal ultrafast quench Sun and Millis (2020); Sefidkhani et al.
+(2024) leading to high domain densities and large values of $d_{0}$.
 Our results are consistent with proposals in which the absorption of energy by the electronic degrees of freedom suppresses the C-CDW order parameter.
@@
 (2023) allowing for the correlations of the H-CDW order parameter to grow Sun and Millis (2020).
-Thus, our results suggest that the minimum of the free energy landscape associated with the H-CDW order parameter already exists at the thermodynamic equilibrium of $1T$-TaS2 and can be accessed with a sufficiently fast adiabatic quench (Figure 3 a).
+Thus, our results and Landau-Ginzburg theory calculations suggest that the minimum of the free energy landscape associated with the H-CDW order parameter already exists at the thermodynamic equilibrium of $1T$-TaS2 and can be accessed with a sufficiently fast adiabatic quench (Figure 4 a).
 Our quenching protocol and the observation of the mixed-CDW phase opens a new pathway towards the development of new efficient functional devices based on $1T$-TaS2 in which H and C-CDW domain engineering can lead to switchable metallic behavior at temperatures above LN2 Mihailovic et al.
@@
 V Acknowledgements
-A.d.l.T acknowledges helpful conversation with Michael Buchold, Yue Cao, Martin Claassen, Greg Fiete, Simon Gerber, and Dante M.
+A.d.l.T acknowledges helpful conversation with Michael Buchold, Yue Cao, Martin Claassen, Simon Gerber, and Dante M.
 Kennes.
@@
 Department of Energy, Office of Science, Office of Basic Energy Sciences, under Award Number DE-SC0021223.
-P.M.V.
-and D.B.
-acknowledge support from NSF DMR 2226097 and the Mason Graduate Division’s Presidential Scholarship Program”.
+P.M.V., S.M.H., and D.B.
+acknowledge support from NSF DMR 2226097 and the Mason Graduate Division’s Presidential Scholarship Program.
 This work is based on research conducted at the Center for High-Energy X-ray Sciences (CHEXS), which is supported by the National Science Foundation (BIO, ENG and MPS Directorates) under award DMR-1829070.
+G.A.F.
+and Y.
+M.
+gratefully acknowlege support from NSF DMR-2114825, and G.A.F.
+from the Alexander von Humboldt Foundation.
 Author Contributions
@@
 W.
-P.
-and J.P.C.R.
+P., and J.P.C.R.
 performed the High Dynamic Range x-ray scattering measurements.
@@
 and P.M.V.
-performed the cryo-Raman experiments.‘ A.d.l.T.
+performed the cryo-Raman experiments.
+Y.
+M.
+and G.A.F.
+performed the Landau-Ginzburg free energy calculations.
+A.d.l.T.
 and.